A right triangle with hypotenuse of length 'a' is rotated about one of its legs to generate a right circular cone. Find the greatest possible volume of such a cone.


1
Expert's answer
2021-03-10T11:39:43-0500

Volume of the cone

"V = \frac13 \pi r^2h" /1/

where

r - radius of the base

h - height


In this problem 'r' and 'h' are the legs of the right triangle with hypotenuse 'a', so

"r^2 + h^2 = a^2" /2/


Express the 'r' in terms of the 'h' using /2/ and substitute it into the expression for volume /1/

"r^2 = a^2 - h^2"


"V = \frac13 \pi (a^2 - h^2)h = \frac13 \pi(a^2h - h^3)" /3/


Find derivative of the volume (assumes that the 'a' is meant to be constant)

"V' = \frac13\pi(a^2 - 3h^2)"


Equate the derivative to zero and calculate 'h'

"\frac13\pi(a^2 - 3h^2) = 0"

"h = \cfrac{a}{\sqrt3}" /4/


Find the second derivative for check is volume has maximum with given 'h'

"V'' = \frac13 \pi(-6h) = -2\pi h"

"V'' < 0" for any positive 'h', so volume has a maximum


Plug obtained 'h' from /4/ into /3/ and find the maximum volume

"V_{max} = \frac13\pi(a^2 - (\frac{a}{\sqrt3})^2)\frac{a}{\sqrt3} = \frac{2}{9\sqrt3}\pi\,a^3"



Answer: greatest possible volume is "\frac{2}{9\sqrt3}\,\pi\,a^3"